Operations Report – January 6th

Crew 238 Operations Report 06-01-2022

Name of person filing report: Simon Werner

Non-nominal systems: 1 CO- Sensor, lower deck, replacement (Atila),
spacesuits, radio, see below.

Notes on non-nominal systems: Spacesuit’s issues/ radio battery, see below

Spirit rover used: No

Hours: (taken today): 149.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: Yes

Hours: 80.2

Beginning charge: 100%

Ending charge: 97%

Currently charging: yes

Curiosity rover used: yes

Hours: 159.8

Beginning charge: 100%

Ending charge: 99%

Currently charging: yes

Perseverance rover used: no

Hours: unknown

Beginning charge: unknown

Ending charge: unknown

Currently charging: unknown

General notes and comments: none

Summary of Hab operations:


Water (static tank): 424

Water (loft tank): 30 gallons, filled to 55 gallons

Water Meter: post pumping 01544750

Static to Loft Pump used – yes

Static tank pipe heater (on or off): on

Static tank heater (On or off) on

Toilet tank emptied: no

Summary of internet: nominal

Summary of suits and radios:.

Spacesuit Nr. 2:

Battery charged to 12.7V, unit back at HAB, switched on for ~1h,
normal airflow, plugged in for re-charge

Spacesuit Nr. 4: Still down, will remove battery tomorrow for charging

One charger seems to be broken, no LED on when connecting to power socket and spacesuit, tested with 3 suits.

All other Spacesuits are nominal and connected to the chargers.

Radio issues: Atila, issue with the one broken battery remains –>
replacement by Outpost possible?

All other radios are nominal and connected to the charger.

Summary of GreenHab operations:

WATER USE: 12 gallons

Heater: On, T=82°F

Supplemental light: No(t) yet

Harvest: Yes, Tomatoes 0.65oz, Mint 0.35oz

Note: inner door handle was loose (screws) –> fixed

Summary of Science Dome operations: Aga investigated the rock samples
collected on our EVA today with the microscope in the middle.

Dual split: Off

Summary of RAM operations:

Used for fixing Spacesuit Nr. 2 (battery installed, re-assembled unit)

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

Found a can of Quakers oats in the RAM this morning – thanks a lot.

Copyright © The Mars Society. All rights reserved. | Main Site